OpenStudy (anonymous):

Suppose that w and t vary inversely and that t = when w = 4. Write a function that models the inverse variation, and find t when w = 9. (1 point)

OpenStudy (anonymous):

please recheck the value of t? it's not stated

OpenStudy (anonymous):

value of t when w = 4,

OpenStudy (anonymous):

while waiting for the value, the working equation you will be using is: \[w = {1(k) \over t}\]

OpenStudy (anonymous):

t=1/5

OpenStudy (anonymous):

just substitute that t to the equation i gave.

OpenStudy (anonymous):

you should be solving the value of k

OpenStudy (anonymous):

after getting the value of k, use again the equation and look for the new value of t.

OpenStudy (anonymous):

I hope this will help

OpenStudy (anonymous):

when w = 4, t = 1/5 w = k/t ----> 4 = k/(1/5) -----> 4 = k/5 therefore k = 20. knowing this substitute the Wnew and the K to the equation solving for Tnew w = k/t -----> t = k/w ----> t = 20/9

OpenStudy (anonymous):

thats not an option though

OpenStudy (anonymous):

haha, im sorry just a the bad of my eye i see the w = 4, as w = 9. Anyway, w=4 and k= 20, so k/w should be 20/4 = 5. sorry.